4RCH4N63L:

A line has a SLOPE = -4 and passes through the POINT (8, -3). Write the equation of the line in slope-intercept form.

heywsp:

whats the question tho-?

4RCH4N63L:

that is the question

SnipingSoup:

do you know the equation form

SmokeyBrown:

The slope of the line is -4, which means that (assuming we are using x-y coordinates as usual), every time x increases by 1, y decreases by 4. We also know that (8, -3) is a point on the line. Knowing the slope, we can work backwards to find the y-intercept (which is the value of y when x is 0). So, if y is -3 when x is 8, y would be (-3 - (8*(-4))) = -3 + 32 = 29 when x is 0. So, now we know the slope (-4) and the y-intercept (29) Can you figure out how to put this information together to write the equation of the line in slope-intercept form?

surjithayer:

eq. of line in slope intercept form is y=mx+c Here m=-4 so eq. of line is y=-4x+c \[\because ~it~passe~through~(8,-3)\] so -3=-4(8)+c c=-3+32=29 Hence eq. of line is y=-4x+29